OpenStudy (anonymous):

the width of a rectangular lot is 8m less than the length. if the area is 105 sq. m, what is the length and the width?

OpenStudy (anonymous):

Start by setting up an equation where W x L = Area = 105 and express W in terms of L. You will have an equation in one variable and it will be straightforward to solve.

OpenStudy (anonymous):

105=WL 15-8=7 15*7=105 L=15 W=7

OpenStudy (anonymous):

@tcarroll010 uhmm.. (w-8) ( w) = 105)?

OpenStudy (anonymous):

ahh ok.